# Detecting eclipsing double white dwarfs with electromagnetic and gravitational waves

### v1

@@ Abstract
Galactic double white dwarfs are predominant sources of gravitational waves in the millihertz frequencies accessible to space-borne gravitational wave detectors. With advances in multi-messenger astronomy, an increasing number of double white dwarf systems will be discovered through both electromagnetic and gravitational wave observations. In this paper, we simulated two populations of double white dwarfs originating from different star formation histories (hereafter referred to as Model 1 and Model 2) using the binary population synthesis method. We predicted the number of double white dwarfs in our Galaxy detectable by TianQin and Laser Interferometer Space Antenna (LISA) individually, as well as through their joint observation. In addition, we performed an analysis to evaluate the accuracy of the parameter estimation using the Fisher information matrix. Furthermore, we predicted the number of eclipsing double white dwarfs detectable by Gaia and LSST. Our study found that over the nominal mission durations, TianQin, LISA, and their joint observation can detect $5\times 10^{3}$ ($1\times 10^{4}$), $1.7\times 10^{4}$ ($3.3\times 10^{4}$), and $1.8\times 10^{4}$ ($3.5\times 10^{4}$) double white dwarfs with signal-to-noise ratios greater than 7 in Model 1 (Model 2), respectively. Gaia and LSST are expected to detect 67 (186) and 273 (554) eclipsing double white dwarfs in Model 1 (Model 2) with orbital period less than 30 hours, respectively. We also found that several dozen eclipsing double white dwarfs can be detected jointly through electromagnetic and gravitational wave observations.
@@
gravitational waves: double white dwarfs.
pubyear: 2022pagerange: Detecting eclipsing double white dwarfs with electromagnetic and gravitational wavesA
@@ 1 Introduction
White dwarfs (WDs) are the final evolutionary stage of main sequence stars with masses less than $8-10~{}{M_{\sun}}$ (Icko Iben et al., 1997; Dobbie et al., 2006). Based on our current understanding of stellar evolution and initial mass distribution, it is estimated that more than 95% of main sequence stars in our Galaxy will evolve into WDs (Baker et al., 2019; Gentile Fusillo et al., 2021a). These stars can no longer produce energy through nuclear fusion reactions, and as a result, they will gradually cool down over time by radiating their stored thermal energy, eventually becoming fainter. The evolution of WDs is reasonably well understood, with further details available in previous reviews (D’Antona & Mazzitelli, 1990; Fontaine et al., 2001; Althaus et al., 2010).
@@
A significant fraction of stars in our Galaxy exist in binary systems (Sana et al., 2012; Duchêne & Kraus, 2013; El-Badry et al., 2021b). Observation shows that the binary fraction is greater than 0.5 for stars with masses $>1~{}{M_{\sun}}$ (see van Haaften et al., 2013, Appendix A). Toonen et al. (2017) simulated a sample of WDs within 20 pc of Earth using a population synthesis method and predicted that optically resolvable double white dwarfs (DWDs) constitute $9-14\%$ of the population. Previous simulations using binary population synthesis have suggested that there are about $10^{8}$ DWDs in our Galaxy (Nelemans et al., 2001b; Yu & Jeffery, 2010; Breivik et al., 2020a). With the advances of multi-messenger astronomy, a wealth of observations on WD binaries will be pivotal in exploring the evolution of binary stars (Postnov & Yungelson, 2014), refining the parameters in binary population synthesis models, such as the efficiency parameter of common envelope evolution (Dewi & Tauris, 2000; Hirai & Mandel, 2022; Chen et al., 2024). In addition, these observations contribute to the exploration of the mechanisms underlying Type Ia supernovae (Toonen et al., 2012) and to the study of the structure of our Galaxy (Rowell & Hambly, 2011; Korol et al., 2019).
@@
Observing WDs is challenging due to their inherently low luminosities. However, the total number of WD candidates has steadily risen with the ongoing progress of large-scale automated surveys and the accumulation of data. The latest WD catalog from Gaia Early Data Release 3 (EDR3) contains about 359,000 high-confidence WD candidates (Gentile Fusillo et al., 2021b). Through a spectroscopic survey focusing on $<0.3~{}{M_{\sun}}$ He-core WDs, the Extremely Low Mass (ELM) survey has discovered 98 detached DWDs (Brown et al., 2020). Follow-up spectroscopy and high-speed photometry observations have confirmed a growing number of DWDs, such as ZTF J153932.16+502738.8 (Burdge et al., 2019), ZTF J0538+1953, ZTF J2029+1534, ZTF J0722+1839, ZTF J1749+0924 (Burdge et al., 2020a), ZTF J2243+5242 (Burdge et al., 2020b), ZTF J1901+5309 (Coughlin et al., 2020), and LAMOST J0140355+392651 (El-Badry et al., 2021a).
Eclipsing binary systems can provide a better opportunity to measure the parameters of both companion stars. We summarized over a dozen known eclipsing DWDs in Table 8 in Appendix A. Based on Gaia’s WD catalog and data from the Zwicky Transient Facility (ZTF), 17 new eclipsing binaries have been identified (Keller et al., 2022), among which the majority are the WD binaries with a main-sequence companion, with only two possibly containing an extremely low-mass WD companion. Utilizing the Gaia EDR3 catalog and ZTF light curve data, Ren et al. (2023) identified a sample of 429 close WD binary candidates through a systematic search for short-period binaries. Wells et al. (2017) quantitatively validated the detection capability of Legacy Survey of Space and Time (LSST; LSST Science Collaboration et al., 2009) for eclipsing binaries by simulating a pseudo-LSST sample of binary light curves and subsequently recovering their periods from the light curves. The results indicate that the periods of $71\%$ binary targets in the sample could be successfully recovered. As the data from Gaia and ZTF continue to accumulate and with the future operation of the 10-year multi-purpose optical survey of LSST, an increasing number of DWDs will be discovered.
@@
On the other hand, in the era of gravitational wave (GW) astronomy, DWDs are one of the most promising sources for proposed space-borne laser interferometers, such as the Laser Interferometer Space Antenna (LISA; Amaro-Seoane et al., 2017), TianQin (Luo et al., 2016), Taiji (Hu & Wu, 2017), ASTROD-GW (Ni, 1998), and gLISA (Tinto et al., 2015). Tens of thousands of DWDs in our Galaxy could be detected by these GW detectors  (Nelemans et al., 2001a; Marsh, 2011; Ruiter et al., 2010; Korol et al., 2017; Lamberts et al., 2019; Huang et al., 2020). Binary systems that have been confirmed by EM observations and are expected to be detected by future GW observations are defined as verification binaries (VBs). Kupfer et al. (2018) collected catalogs of known VBs and identified 11 VBs with a signal-to-noise ratios (SNRs) greater than 20, two additional VBs with SNRs of at least 5, and three additional VBs projected to archive SNRs of approximately 5 following a 4-year mission duration of LISA. Burdge et al. (2020a) conducted a systematic search of the ZTF data and identified 15 binary systems with orbital periods ranging from 6.91 to 56.35 minutes. Among these, 5 sources have SNR greater than 5 for a 4-year mission duration of LISA. For joint GW observations, it was found that the spatial localization and characterization of binaries can be improved by combining data from multiple detectors, including LISA, TianQin, and Taiji (Gong et al., 2021; Wu & Li, 2023).
@@
While more than one hundred DWD systems have been discovered so far by electromagnetic (EM) observations, there is a noticeable selection bias. This bias arises from the fact that DWDs with higher luminosities and closer distances to the observer are more likely to be detected by EM observations. However, to study the population characteristics of DWDs in our Galaxy, we need a more comprehensive sample. Binary population synthesis (BPS) is the conventional method for studying the evolution of binary populations and has wide applications in astrophysics. To date, a variety of BPS codes have been developed, such as SeBa (Portegies Zwart & Verbunt, 1996; Nelemans et al., 2001a), Yunnan Model (Han, 1998; Han et al., 2002, 2003; Zhang et al., 2002, 2004), BSE (Hurley et al., 2002), COSMIC (Breivik et al., 2020b), and COMPAS (Riley et al., 2022). Binary population synthesis produces statistical samples of certain types of stars for subsequent astrophysical investigations. For example, it has been used to explore the number of detectable DWDs in our Galaxy for LISA (Toonen et al., 2017; Korol et al., 2020). Huang et al. (2020) investigated the ability of TianQin to detect DWDs in our Galaxy using a synthetic catalog from  Toonen et al. (2012, 2017), which shows that about $8.7\times 10^{3}$ DWDs can be detected over the 5-year nominal mission duration of TianQin. Finch et al. (2023) found that prior EM-derived knowledge of VBs can improve the accuracy of GW parameter estimation in the Bayesian framework. The formation and evolution of ultra-compact X-ray binaries (UCXBs) in the Galactic bulge has been explored using SeBa simulations by van Haaften et al. (2013). Renzo et al. (2021) investigated the potential of LISA to detect of GWs from inspiraling binaries in common envelope phases using COSMIC. Lau et al. (2020) used a double neutron star (DNS) population generated by COMPAS to estimate the properties of DNS systems with cumulative SNRs greater than 8 during the 4-year LISA mission. Using a modified BSE code, Shao & Li (2021) predicted that LISA could detect dozens of GW sources with stellar-mass black holes in our Galaxy.
In this paper, we used the Compact Object Synthesis and Monte Carlo Investigation Code (COSMIC; Breivik et al., 2020b), a publicly available Python-based population synthesis software [^1] https://cosmic-popsynth.github.io/docs/stable/, to generate comprehensive samples of DWDs in our Galaxy. We simulated two populations of DWDs based on different star formation histories (hereafter referred to as Model 1 and Model 2). We find that the number of detectable DWDs with GWs or EM waves is 2-3 times higher in Model 2 than in Model 1. We simulated the light curves of DWDs in the Gaia $G$-band and LSST $(ugrizy)$ bands using the PHysics Of Eclipsing BinariEs (PHOEBE; Prša, 2018). Our simulations show that the LSST-$r$ band performs best compared to the other bands. For future multi-messenger observations of DWDs, we predict the differences in the parameter distributions of DWDs detectable with GW and EM observations. We find that EM observations are more likely to detect DWDs with longer periods and closer distances compared to GWs.
@@
The structure of this paper is as follows. In Section 2, we introduce the population synthesis method used to generate DWD populations in our Galaxy. In Section 3, we simulate the light curves of eclipsing DWDs and investigate their detectability in the photometric surveys of Gaia and LSST. In Section 4, we estimate the number of DWDs that are detectable by LISA, TianQin, and their joint observations. In addition, we evaluate the accuracy of the parameter estimation for these DWDs. In Section 5, we discuss the results of the joint observations of DWDs with GW and EM observations. The paper is concluded in Section 6.
@@ 2.1 Initial parameters
We initiate a population of zero-age main-sequence (ZAMS) stars by employing the probability distribution functions for their initial parameters. Subsequently, we evolve this initial population using a specific evolutionary model which leads to the formation of a sample of DWDs. The initial values for the parameters of the binaries are derived by sampling the distribution functions related to the initial metallicity ($Z$), primary mass ($m$), binary mass ratio ($q$), orbital separation ($a$), eccentricity ($e$), and orbital inclination angle ($\iota$). COSMIC has several built-in options for the distribution functions of the initial parameters, such as the initial mass functions: ‘Salpeter55’ (Salpeter, 1955), ‘Kroupa93’ (Kroupa et al., 1993), ‘Kroupa01’ (Kroupa, 2001), or user-defined power-law exponents; the initial orbital separation distribution functions: log-uniform (Abt, 1983), ‘Sana2012’ (Dominik et al., 2012), ‘Renzo2019’ (Renzo et al., 2019); the initial eccentricity distributions: the thermal distribution (Heggie, 1975), the uniform distribution (Geller et al., 2019). In our simulations, the initial mass function follows Kroupa et al. (1993) with a mass range of $0.08-150~{}M_{\sun}$, the binary mass ratio $q$ adopts the uniform distribution between $[0,1]$, the orbital separation (in unit of solar radius) is sampled from the log-uniform distribution with $\log(a)$ between $[0,6]$, and the eccentricity adopts the thermal distribution with $e$ between $[0,1]$. We use a constant binary fraction of 0.5 and a constant metallicity of 0.02.
To investigate the impact of star formation history on binary system evolution, we simulate two populations of DWDs that share the same initial parameters described in the previous paragraph except for their star formation histories. In Model 1, we adopt a burst of star formation 13.7 Gyr ago (Yu & Jeffery, 2010). In Model 2, we follow the star formation history in Wagg et al. (2022), which assumes a constant star formation rate over 10 Gyr for the thin disk, a 1 Gyr burst of star formation 11 Gyr ago for the thick disk, and a 1 Gyr burst of 10 Gyr ago for the bulge.
Table 1: Science requirements for Gaia and LSST. The referenced parameters are from Gaia Collaboration et al. (2016a) and Ivezić et al. (2019), respectively.
Gaia LSST
Survey Area whole sky 20000 ${\rm deg}^{2}$
Single Visit Limiting Magnitude 21 $u$: 23.9; $g$: 25; $r$: 24.7;
$i$: 24; $z$: 23.3: $y$: 22.1
Total visits per sky patch 70 $u$: 56; $g$: 80; $r$: 184;
$i$: 184; $z$: 160: $y$: 160
Wavelength coverage 330–1050 nm 320–1050 nm ($ugrizy$)
Nominal mission lifetime 5yr 10 yr
@@ 2.2 Galactic structure
The position of the DWD in our Galaxy is essential for calculating the flux of the light curve (Sec. 3.1) and the signal-to-noise ratio (SNR) of the GW signals (Sec. 4.1). Here, we assume that the structure of our Galaxy is axisymmetric, and assign a position to each pair of DWDs by sampling the distribution of the mass densities in different Galactic structures. Previous work usually assumed an exponential squared bulge and a squared hyperbolic secant plus an exponential disk or a double exponential disk (Nelemans et al., 2004; Yu & Jeffery, 2010; Korol et al., 2019). A comprehensive summary of the Galactic models widely employed in the literature is provided in Nissanke et al. (2012). In this work, we adopt the density profile for the bulge described in McMillan (2011):
$\rho_{\mathrm{b}}(R,z)=\frac{\rho_{\mathrm{b},0}}{\left(1+r^{\prime}/r_{0}\right)^{\alpha}}\exp\left[-\left(r^{\prime}/r_{\mathrm{cut}}\right)^{2}\right]~{}~{}{M_{\sun}\,\rm{kpc^{-3}}}\,,$ (2)
where the cylindrical polar coordinate $R$ is the radial distance and $z$ is the height. $r^{\prime}=\sqrt{R^{2}+(z/q_{1})^{2}}$, $q_{1}=0.5$, $\alpha=1.8$, $r_{0}=0.075\,\mathrm{kpc}$, $r_{\mathrm{cut}}=2.1\,\mathrm{kpc}$, and $\rho_{\mathrm{b},0}=9.93\times 10^{10}\,M_{\sun}\,\rm{kpc^{-3}}$.
The Galaxy’s disk is composed of a thin and a thick part, each with a double-exponential density profile with different scale lengths and scale heights,
$\rho_{\mathrm{d}}(R,z)=\frac{\Sigma_{\mathrm{d},0}}{2z_{\mathrm{d}}}\exp\left(-\frac{|z|}{z_{\mathrm{d}}}-\frac{R}{R_{\mathrm{d}}}\right){M_{\sun}}\,\mathrm{kpc^{-3}}\,.$ (3)
For the thin disk, the scale height is $z_{\mathrm{d}}=0.3\,\mathrm{kpc}$, the scale length is $R_{\mathrm{d}}=2.9\,\mathrm{kpc}$, and the central surface density is $\Sigma_{\mathrm{d,0}}=816.6\,{M_{\sun}\,\rm{pc^{-2}}}$. For the thick disk, the scale height is $z_{\mathrm{d}}=0.9\,\mathrm{kpc}$, the scale length is $R_{\mathrm{d}}=3.31\,\mathrm{kpc}$, and the central surface density is $\Sigma_{\mathrm{d,0}}=209.5\,{M_{\sun}\,\mathrm{pc}^{-2}}$. For the Sun-Galactic center distance, we adopt $R_{0}=8.178\,\mathrm{kpc}$ (GRAVITY Collaboration et al., 2019).

### v2

@@ Abstract
Galactic double white dwarfs are predominant sources of gravitational waves in the millihertz frequencies accessible to space-borne gravitational wave detectors. With advances in multi-messenger astronomy, an increasing number of double white dwarf systems will be discovered through both electromagnetic and gravitational wave observations. In this paper, we simulated two populations of double white dwarfs originating from different star formation histories (hereafter referred to as Model 1 and Model 2) using the binary population synthesis method. We predicted the number of double white dwarfs in our Galaxy detectable by TianQin and Laser Interferometer Space Antenna (LISA) individually, as well as through their joint observation. In addition, we performed an analysis to evaluate the accuracy of the parameter estimation using the Fisher information matrix. Furthermore, we predicted the number of detached eclipsing double white dwarfs detectable by Gaia and the Vera C. Rubin Observatory (VRO). Our study found that over the nominal mission durations, TianQin, LISA, and their joint observation can detect at least five thousand and potentially several tens of thousands of double white dwarfs with signal-to-noise ratios greater than 7. Gaia and VRO are expected to detect at least several dozen and up to several hundred eclipsing double white dwarfs with orbital periods less than 30 hours. We also found that several dozen eclipsing double white dwarfs can be detected jointly through electromagnetic and gravitational wave observations.
@@
gravitational waves – binaries: close – binaries: eclipsing – white dwarfs.
pubyear: 2022pagerange: Detecting eclipsing double white dwarfs with electromagnetic and gravitational wavesDetecting eclipsing double white dwarfs with electromagnetic and gravitational waves
@@ 1 Introduction
White dwarfs (WDs) are the final evolutionary stage of main sequence stars with masses less than $8-10~{}{M_{\sun}}$ (Icko Iben et al., 1997; Dobbie et al., 2006). Based on our current understanding of stellar evolution and initial mass function, it is estimated that more than 95% of main sequence stars in our Galaxy will evolve into WDs (Baker et al., 2019; Gentile Fusillo et al., 2021a). These stars can no longer produce energy through nuclear fusion reactions, and as a result, they will gradually cool down over time by radiating their stored thermal energy, eventually becoming fainter. The evolution of WDs is reasonably well understood, with further details available in previous reviews (D’Antona & Mazzitelli, 1990; Fontaine et al., 2001; Althaus et al., 2010). For a discussion of recent advances and challenges in observations and theoretical modeling, see (Saumon et al., 2022; Tremblay et al., 2024; Bédard, 2024).
@@
A significant fraction of stars in our Galaxy reside in binary systems (Sana et al., 2012; Duchêne & Kraus, 2013a; El-Badry et al., 2021b). Observation shows that the binary fraction is greater than 0.5 for stars with masses $>1~{}{M_{\sun}}$ (see van Haaften et al., 2013, Appendix A). Toonen et al. (2017) simulated a sample of WDs within 20 pc of Earth using a population synthesis method and predicted that optically resolvable (angular separation of orbit larger than a critical value) double white dwarfs (DWDs) constitute $9-14\%$ of the population. Previous simulations using binary population synthesis have suggested that there are about $10^{8}$ DWDs in our Galaxy (Nelemans et al., 2001b; Yu & Jeffery, 2010; Breivik et al., 2020a). With the advances of multi-messenger astronomy, a wealth of observations on WD binaries will be pivotal in exploring the evolution of binary stars (Postnov & Yungelson, 2014), refining the parameters in binary population synthesis models, such as the efficiency parameter of common envelope evolution (Dewi & Tauris, 2000; Hirai & Mandel, 2022; Chen et al., 2024). In addition, these observations contribute to the exploration of the mechanisms underlying Type Ia supernovae (Toonen et al., 2012) and to the study of the structure of our Galaxy (Rowell & Hambly, 2011; Korol et al., 2019).
@@
Observing WDs is challenging due to their inherently low luminosities. However, the total number of WD candidates has steadily risen with the ongoing progress of large-scale automated surveys and the accumulation of data. The latest WD catalog from Gaia Early Data Release 3 (EDR3) contains about 359,000 high-confidence WD candidates (Gentile Fusillo et al., 2021b). Through a spectroscopic survey focusing on $<0.3~{}{M_{\sun}}$ He-core WDs, the Extremely Low Mass (ELM) survey has discovered 98 detached DWDs (Brown et al., 2020), and more recently an additional 34 ELM WD binaries have been discovered in the southern sky (Kosakowski et al., 2020; Kosakowski et al., 2023).Follow-up spectroscopy and high-speed photometry observations have confirmed a growing number of DWDs (Burdge et al., 2019, 2020a, 2020b; Coughlin et al., 2020; El-Badry et al., 2021a).
[…]
@@
On the other hand, in the era of gravitational wave (GW) astronomy, DWDs are one of the most promising sources for proposed space-borne laser interferometers, such as the Laser Interferometer Space Antenna (LISA; Amaro-Seoane et al., 2017), TianQin (Luo et al., 2016), Taiji (Hu & Wu, 2017), ASTROD-GW (Ni, 1998), and gLISA (Tinto et al., 2015). Tens of thousands of DWDs in our Galaxy could be detected by these GW detectors  (Nelemans et al., 2001a; Marsh, 2011; Ruiter et al., 2010; Korol et al., 2017; Lamberts et al., 2019; Huang et al., 2020). Binary systems that have been confirmed by EM observations and are expected to be detected by future GW observations are defined as verification binaries (VBs). To date, several tens of VBs have been discovered and their properties have been well characterized (Kupfer et al., 2018, 2024; Burdge et al., 2020a; Munday et al., 2024).
@@
While more than one hundred DWD systems have been discovered so far by electromagnetic (EM) observations, there is a noticeable selection bias. This bias arises from the fact that DWDs with higher luminosities and closer distances to the observer are more likely to be detected by EM observations. However, to study the population characteristics of DWDs in our Galaxy, we need a more comprehensive sample. Binary population synthesis (BPS) is the conventional method for studying the evolution of binary populations and has wide applications in astrophysics. To date, a variety of BPS codes have been developed, such as SeBa (Portegies Zwart & Verbunt, 1996; Nelemans et al., 2001a), Yunnan Model (Han, 1998; Han et al., 2002, 2003; Zhang et al., 2002, 2004), BSE (Hurley et al., 2002), COSMIC (Breivik et al., 2020b), and COMPAS (Riley et al., 2022). Binary population synthesis produces statistical samples of certain types of stars for subsequent astrophysical investigations. For example, it has been used to explore the number of detectable DWDs in our Galaxy for LISA (Toonen et al., 2017; Korol et al., 2020). Huang et al. (2020) investigated the ability of TianQin to detect DWDs in our Galaxy using a synthetic catalog from  Toonen et al. (2012, 2017), which shows that about $8.7\times 10^{3}$ DWDs can be detected over the 5-year nominal mission duration of TianQin. The formation and evolution of ultra-compact X-ray binaries (UCXBs) in the Galactic bulge has been explored using SeBa simulations by van Haaften et al. (2013). Renzo et al. (2021) investigated the potential of LISA to detect of GWs from inspiraling binaries in common envelope phases using COSMIC. Lau et al. (2020) used a double neutron star (DNS) population generated by COMPAS to estimate the properties of DNS systems with cumulative SNRs greater than 8 during the 4-year LISA mission. Using a modified BSE code, Shao & Li (2021) predicted that LISA could detect dozens of GW sources with stellar-mass black holes in our Galaxy.
@@
The structure of this paper is as follows. In Section 2, we introduce the population synthesis method used to generate DWD populations in our Galaxy. In Section 3, we simulate the light curves of eclipsing DWDs and investigate their detectability in the photometric surveys of Gaia and VRO. In Section 4, we estimate the number of DWDs that are detectable by LISA, TianQin, and their joint observations. In addition, we evaluate the accuracy of the parameter estimation for these DWDs. In Section 5, we discuss the results of the joint observations of DWDs with GW and EM observations. The paper is concluded in Section 6.
@@ 2.1 Initial parameters
We initiate a population of zero-age main-sequence (ZAMS) stars by employing the probability distribution functions for their initial parameters. Subsequently, we evolve this initial population using a specific evolutionary model which leads to the formation of a sample of DWDs. The initial values for the parameters of the binaries are derived by sampling the distribution functions related to the initial metallicity ($Z$), primary mass ($m$), binary mass ratio ($q$), orbital separation ($a$), eccentricity ($e$), and orbital inclination angle ($\iota$). In our simulations, the initial mass function follows Kroupa et al. (1993) with a mass range of $0.08-150~{}M_{\sun}$, which provides a universally applicable model based on Galactic field data, reflects observed power-law index changes at critical masses ($0.08~{}M_{\sun}$ and $0.5~{}M_{\sun}$), and effectively quantifies observational uncertainties such as Poisson noise and dynamical evolution (Kroupa, 2001); The binary mass ratio $q$ adopts the uniform distribution between $[0,1]$ and this flat mass ratio distribution is a good first-order approximation when compared to observations (Duchêne & Kraus, 2013b); The orbital separation (in unit of solar radius) is sampled from the log-uniform distribution with $\log(a)$ between $[0,6]$ (Abt, 1983), which is one of the most commonly used assumptions for the orbital semimajor axis distribution and is consistent with observations for solar mass and intermediate mass stars (Duchêne & Kraus, 2013b); and the eccentricity adopts the thermal distribution with $e$ between $[0,1]$(Heggie, 1975), which reflects the expectation from energy equipartition. In addition, because DWDs quickly circularize orbits in their early evolution, the initial eccentricity distribution has a limited effect on the simulation results. We use a constant binary fraction of 0.5 and a constant metallicity of 0.02. Our initial parameters are the same as those in Breivik et al. (2020a), but differ from the mass range ($0.95-10~{}M_{\sun}$) for the initial mass function used in Korol et al. (2019).
To investigate the impact of star formation history on binary system evolution, we simulate two populations of DWDs that share the same initial parameters described in the previous paragraph except for their star formation histories. In Model 1, we adopt a burst of star formation 13.7 Gyr ago (Yu & Jeffery, 2010), which can serve as a useful initial choice. Other stellar populations can then be obtained from this choice through convolution with specific star formation histories. In Model 2, we follow the star formation history in Wagg et al. (2022), which assumes a constant star formation rate over 10 Gyr for the thin disk, a 1 Gyr burst of star formation 11 Gyr ago for the thick disk, and a 1 Gyr burst of 10 Gyr ago for the bulge. This can be taken as an average of the star formation rate over certain cosmological epoch and is usually appropriate for low-redshift spiral galaxies (Boquien et al., 2014).
@@ 2.2 Galactic structure
The position of the DWD in our Galaxy is essential for calculating the flux of the light curve (Sec. 3.1) and the signal-to-noise ratio (SNR) of the GW signals (Sec. 4.1). Here, we assume that the structure of our Galaxy is axisymmetric, and assign a position to each pair of DWDs by sampling the distribution of the mass densities in different Galactic structures. Previous work usually assumed an exponential squared bulge and a squared hyperbolic secant plus an exponential disk or a double exponential disk (Nelemans et al., 2004; Yu & Jeffery, 2010; Korol et al., 2019). A comprehensive summary of the Galactic models widely employed in the literature is provided in Nissanke et al. (2012). In this work, we adopt the density profile and the values of the associated parameters for the bulge, thin disk, and thick disk described in McMillan (2011), except for the Sun-Galactic center distance, for which we adopt $R_{0}=8.178\,\mathrm{kpc}$ (GRAVITY Collaboration et al., 2019).